Silver sulfide and hydrogen gas combine to form silver and hydrogen sulfide. 3.0 g of H2 and 100 g of Ag2S are used to produce silver. The equation for the reaction is:

 Ag2S + H2 -> 2Ag + H2S

Calculate the theoretical yield of silver? Your answer needs to explain what is the limiting reagent.


1
Expert's answer
2021-06-11T01:57:58-0400

According to the reaction one mole of hydrogen reacts with one mole of silver sulfide. The number of moles of reactants used in the reaction are:

n(H2) = m(H2) / Mr(H2) = 3.0 g / 2 g/mol = 1.5 mol

n(Ag2S) = m(Ag2S) / Mr(Ag2S) = 100 g / 248 g/mol = 0.4 mol

Since the reaction occurs between 0.4 mol of Ag2S and 1.5 mol of H2, Ag2S is a limiting reactant. From here, the theoretical yield of silver equals:

m(Ag) = m(Ag2S) × Mr(Ag) × 2 / Mr(Ag2S) = 100 g × 108 g/mol × 2 / 248 g/mol = 87.1 g


Answer: 87.1 g